OpenStudy (anonymous):

w^2-2w+1=0 solve by completing the square

OpenStudy (saifoo.khan):

w=1

myininaya (myininaya):

\[(w-1)^2=0\]

OpenStudy (anonymous):

\[w^2-2w=-1\] \[(w-1)^2=-1+1=0\] \[(w-1)^2=0\] \[w-1=0\] \[w=1\]

OpenStudy (anonymous):

has to have two answers it a second degree equation

OpenStudy (saifoo.khan):

No, there will be just one. this is because the curve only intersects the x-axis once,

myininaya (myininaya):

1 multiplicity 2

OpenStudy (anonymous):

okay i wrote problem wrong it is w^2-2w-63

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

sorry

myininaya (myininaya):

-63=-9*7 -2=-9+7

OpenStudy (saifoo.khan):

\[(w-9)(w+7)=0\] w = 9, w = -7

myininaya (myininaya):

\[(w-9)(w+7)=0\]

OpenStudy (anonymous):

okay i just had my signs wrong and now i see where my mistake was thank you

OpenStudy (saifoo.khan):

you welcome.

OpenStudy (anonymous):

okay how about w^2+2w=4 solve by factoring or completeing the square

OpenStudy (anonymous):

i move the 4 over so it wold become w^2+2w-4=0

OpenStudy (saifoo.khan):

\[\Large -1 \sqrt5, \sqrt5 -1\]

OpenStudy (anonymous):

i would not move it over. i would write \[w^2+2x=4\] \[(w+1)^2=4+1=5\] \[w+1=\pm\sqrt{5}\] \[w=-1\pm\sqrt{5}\]

OpenStudy (anonymous):

okay i got it i was trying to factor...but i see i can't

myininaya (myininaya):

w and x?

OpenStudy (saifoo.khan):

this cannot be factored.

myininaya (myininaya):

satellite how did you solve your equation? oh wait does w=x?

myininaya (myininaya):

lol

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

okay so how about 8s-11=(2-6)(s-4)

myininaya (myininaya):

8+4=16?

OpenStudy (saifoo.khan):

Dang!

myininaya (myininaya):

lol

OpenStudy (anonymous):

i mistyped the problems ready 8s-11=(s-6)(s-4) i am horriable typer

OpenStudy (saifoo.khan):

LOL.

myininaya (myininaya):

multiply the left hand side then put everything on one side and combine like terms

OpenStudy (saifoo.khan):

multiply left side or right side?

OpenStudy (anonymous):

so i got 8s-11=s^2-10s+24

OpenStudy (anonymous):

is that right?

myininaya (myininaya):

i don't know my left from right do whatever side that requires multiplication

OpenStudy (saifoo.khan):

\[\huge \checkmark\]

OpenStudy (saifoo.khan):

LOL.

myininaya (myininaya):

yes thats right cas

OpenStudy (anonymous):

okay so then combine the sides and i got s^2-18s_35 is that correct

myininaya (myininaya):

goodnight guys i have to read a paper good luck cas and i will be watching for mistakes saifoo :)

OpenStudy (saifoo.khan):

Good night!! :D

OpenStudy (anonymous):

thanks and goodnight

OpenStudy (saifoo.khan):

\[-s^2+18 s-35 = 0 \]

OpenStudy (anonymous):

okay so what is the next step now

OpenStudy (saifoo.khan):

\[s = 9-\sqrt{46} \] \[s = 9+\sqrt{46}\]

OpenStudy (anonymous):

how do you get that?

OpenStudy (saifoo.khan):

Quadratic formula!

OpenStudy (anonymous):

okay thanks

OpenStudy (saifoo.khan):

no problem,

OpenStudy (anonymous):

how about 8 plus or minus square root of -36 over 2

OpenStudy (saifoo.khan):

\[\frac{8 \pm \sqrt{-36}}{2}\]Like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[4+3i, 4-3i\]

OpenStudy (anonymous):

all right wise guys w = x!

OpenStudy (anonymous):

so how do it can you do it step by step for me trying to really understand it

OpenStudy (saifoo.khan):

\[\frac{8+ 6i}{2}, \frac{8-6i}{2}\] \[\frac{2(4+3i)}{2},\frac{2(4-3i)}{2}\] Both 2 cancels out, finally, \[4+3i, 4-3i\]

OpenStudy (anonymous):

okay thank you. i am finally starting to feel most comfortable

OpenStudy (saifoo.khan):

You welcome.

OpenStudy (anonymous):

so z^2-2z-191 1. \[-2\pm \sqrt{4^{2}}-4(1)(-191) \over 2\]

OpenStudy (saifoo.khan):

\[-2\pm \sqrt{4^{2}-4(1)(-191)} \over 2\]

OpenStudy (anonymous):

okay so then it become \[2\pm \sqrt{4}-(-764)\] over 2

OpenStudy (saifoo.khan):

No, the square root is whole thing.

OpenStudy (anonymous):

yes yours it correct still learning how to you it

OpenStudy (anonymous):

use not you

OpenStudy (saifoo.khan):

Ohh, np

OpenStudy (anonymous):

oksy so then it is \[2\pm \sqrt{-768} \over 2 \]

OpenStudy (saifoo.khan):

\[2\pm \sqrt{780} \over 2\]

OpenStudy (anonymous):

then \[2\pm \sqrt{256 \times3} \over 2 \]

OpenStudy (anonymous):

768 not 780

OpenStudy (saifoo.khan):

\[\sqrt{780} = 2 \sqrt{195}\]

OpenStudy (anonymous):

then \[2\pm16 \sqrt{3} \over 2\]

OpenStudy (saifoo.khan):

it is 780.

OpenStudy (anonymous):

are you solving \[z^2-2z-191=0 \]?

OpenStudy (anonymous):

not sorry trying to go to fast 768 yes

OpenStudy (saifoo.khan):

16+764 ===> 780

OpenStudy (anonymous):

\[z-2z=191\] \[(z-1)^2+191+1=192\] \[z-1=\pm\sqrt{192}=8\sqrt{3}\] \[z=1\pm8\sqrt{3}\]

OpenStudy (anonymous):

if the "middle term" has an even coefficient it is much easier to complete the square. that is because you will not have a denominator that the quadratic equation forces on you

OpenStudy (anonymous):

in any case cassy you have it because as soon as you divide your answer by the 2 in the denominator you will get \[1\pm8\sqrt{3}\]

OpenStudy (anonymous):

okay that makes since thanks

OpenStudy (saifoo.khan):

\[\Huge \checkmark\]

OpenStudy (anonymous):

hello saifooo!

OpenStudy (saifoo.khan):

Hey Sat. xD Openstudy is so low these days! :(

OpenStudy (anonymous):

4w^2-36w+78=0

OpenStudy (saifoo.khan):

\[\frac12(9-\sqrt3), \frac12(9+\sqrt3)\]

OpenStudy (anonymous):

so solving by quadratic formula right not by factor

OpenStudy (saifoo.khan):

Yes, that's easy. and it dosn't wastes our time. secondly, it has less errors!

OpenStudy (anonymous):

okay great thanks i think i may have a few more questions

OpenStudy (saifoo.khan):

Sure..

OpenStudy (anonymous):

4t^2-18t+39=0

OpenStudy (anonymous):

do i do quad or factor

OpenStudy (saifoo.khan):

Quad.

OpenStudy (anonymous):

and if i do quad whne i get to \[18\pm \sqrt{-300} \over 8\] iam lost

OpenStudy (saifoo.khan):

\[\huge \checkmark \]

OpenStudy (anonymous):

now what do i do do i factor out 300 to what?

OpenStudy (saifoo.khan):

\[\sqrt{-300}=\sqrt{-100 \times 3}=-10i \sqrt3\]

OpenStudy (anonymous):

then the \[18pm10i \sqrt{3} \over 8\] is that te moreh answer or can i simply

OpenStudy (anonymous):

is that the answer or can i simply fly

OpenStudy (saifoo.khan):

we can simplify by taking 2 common